On January 1,2020, Drenz Co. acquired 25% interest in the ordinary shares of Josiah Inc. for ₱3,000,000 which reflected book value as of that date. Josiah Company reported net income and paid dividends for 2020 and 2021 as follows:

2020 2021

Net income ₱1,000,000 ₱1,500,000

Dividend declared 400,000 700,000

Questions: 

Case No. 1: On November 20,2020, Drenz Company sold inventory costing ₱60,000 to Josiah Co. for ₱110,000, 60% of which was still unsold on December 31,2020.

1. Investment income 2020

   2. Investment 2021

Case No. 2: On November 20,2020, Josiah company sold inventory costing ₱70,000 to Drenz Co. for 120,000, 60% of which was still unsold on December 31,2020.

3.Investment income in 2020

4. Investment income 2021



1
Expert's answer
2021-06-17T12:57:57-0400

Balance increasing:

"2,000,000-800,000+4,500,000-1,600,000-\frac{2,500,000}{5}*2=3,100,000"

Balance book values on On December 31, 2022:

"10,000,000+3,100,000=13,100,000"

Drenz Co. has 30% of Josiah Inc.

the carrying amount of the investment as of December 31,2022:

"13,100,000*0.3=3,930,000"